# Understanding the Training and Generalization of Pretrained Transformer for Sequential Decision Making

### v1

@@ Abstract
In this paper, we consider the supervised pretrained transformer for a class of sequential decision-making problems. The class of considered problems is a subset of the general formulation of reinforcement learning in that there is no transition probability matrix, and the class of problems covers bandits, dynamic pricing, and newsvendor problems as special cases. Such a structure enables the use of optimal actions/decisions in the pretraining phase, and the usage also provides new insights for the training and generalization of the pretrained transformer. We first note that the training of the transformer model can be viewed as a performative prediction problem, and the existing methods and theories largely ignore or cannot resolve the arisen out-of-distribution issue. We propose a natural solution that includes the transformer-generated action sequences in the training procedure, and it enjoys better properties both numerically and theoretically. The availability of the optimal actions in the considered tasks also allows us to analyze the properties of the pretrained transformer as an algorithm and explains why it may lack exploration and how this can be automatically resolved. Numerically, we categorize the advantages of the pretrained transformer over the structured algorithms such as UCB and Thompson sampling into three cases: (i) it better utilizes the prior knowledge in the pretraining data; (ii) it can elegantly handle the misspecification issue suffered by the structured algorithms; (iii) for short time horizon such as $T\leq 50$, it behaves more greedy and enjoys much better regret than the structured algorithms which are designed for asymptotic optimality.
@@ 1 Introduction
In recent years, transformer-based models (Vaswani et al., 2017) have achieved great success in a wide range of tasks such as natural language processing (Radford et al., 2019; Bubeck et al., 2023), computer vision (Dosovitskiy et al., 2020), and also reinforcement learning (RL) (Chen et al., 2021). In particular, an offline RL problem can be framed as a sequence prediction problem (Ding et al., 2019) that predicts the optimal/near-optimal/human action based on the observed history. Chen et al. (2021); Janner et al. (2021) pioneer this approach with the Decision Transformer (DT), which leverages the auto-regressive nature of transformers to maximize the likelihood of trajectories in offline datasets. In this paper, we apply the idea of supervised pretrained transformer on a subclass of reinforcement learning problems, which we call sequential decision-making problems, that has no transition probability matrix compared to a general RL problem. This simplified problem structure brings two benefits: (i) the optimal action is either in closed form or efficiently solvable and thus we can use the optimal action as the prediction target in the training data and (ii) the setting is more aligned with the existing study (Xie et al., 2021; Garg et al., 2022) on the in-context learning ability of the transformer which gives more insight to the working machinery of transformer on decision-making tasks. Moreover, the class of sequential decision-making problems studied in this paper, covers stochastic and linear bandit problems (Lattimore and Szepesvári, 2020), dynamic pricing (Den Boer, 2015; Cohen et al., 2020), and newsvendor problem (Levi et al., 2015; Besbes and Muharremoglu, 2013) as special cases. Therefore, as a side product, our paper extends pretrained transformers to the application contexts of business, economics, and operations research which are less studied by the existing works.
@@
- The training procedure: In generating the training data for the pretrained transformer, it needs to generate both the actions in the history (which are the features) and the optimal/near-optimal action to predict (which are the target variable). We point out that when these two streams of actions are generated by different algorithms/methods, it will result in an out-of-distribution (OOD) issue largely ignored or unsolved by the existing literature. Zheng et al. (2022) show that when both streams are generated by one structured algorithm (say UCB or Thompson sampling), the pretraining of the transformer will result in an imitation learning procedure that imitates the structured algorithm at best. If one desires to obtain a better algorithm than the structured algorithm, like using the optimal action (Lee et al., 2024) as the target to predict in our case, this will cause the OOD issue (reflected by an $\Theta(\exp T)$ factor in the regret bound of Zheng et al. (2022)). To resolve this matter, we propose an algorithmic solution by injecting the transformer-generated action sequence into the pretraining data, and this also draws a connection with the problem of performative prediction (Perdomo et al., 2020).
- Transformer as decision-making algorithms: The transformer is trained to minimize the empirical risk as the standard supervised learning, and hence the pretrained transformer will approximate the Bayes-optimal predictor for the supervised loss as best. Yet, when used for decision-making problems, the transformer works as a decision-making algorithm. We investigate the gap between the prediction and the decision-making by interpreting the transformer from a Bayesian posterior perspective (Xie et al., 2021; Zhang et al., 2023b; Bai et al., 2024), and in particular, for the continuous action space. The investigation explains (i) why the transformer model exhibits a lack of exploration and (ii) how it achieves a paramount or better performance than the structured algorithms.
@@
- When the transformer is better: While there have been many structured algorithms for each of the decision-making problems that the transformer model is applied to, the question is whether and when the transformer can achieve a better performance than these algorithms. We use numerical experiments to attribute the advantage of the pretrained transformer into three scenarios: (i) it better utilizes the prior knowledge in the pretraining data; (ii) it can elegantly handle the misspecification issue suffered by the structured algorithms; (iii) for short time horizon such as $T\leq 50$ or $100$, it behaves more greedy and enjoys much better regret than the structured algorithms which are designed for asymptotic optimality.
@@ 2 Problem Setup
In this section, we introduce a general formulation for sequential decision-making and describe the setup of supervised pretraining. An agent (decision maker) makes decisions in a sequential environment. The full procedure is described by the sequence
@@
where $\theta$ encapsulates all the parameters of the function. That is, the agent learns from the past interactions with the underlying environment – tuples of $(X_{t},a_{t},o_{t})$’s to optimize future decisions, and this process that transfers the past knowledge to the future is described by the function $\texttt{TF}_{\theta}$. The function $\texttt{TF}_{\theta}$ can take as inputs a variable-length sequence ($H_{t}$ can be of variable length), which is the case for the generative pretrained transformers (Radford et al., 2019).
@@ 2.1 Environment and performance metrics
2.2 Supervised pretraining
In the classic study of these sequential decision-making problems, the common paradigm is to assume some structure for the underlying environment and/or the reward function, and to design algorithms accordingly. These algorithms are often known as online or reinforcement learning algorithms and are usually hard to combine with prior knowledge such as pretraining data. Comparatively, the supervised pretraining formulates the decision-making problem as a supervised learning prediction task that predicts the optimal or a near-optimal action, and thus it can utilize the vast availability of pretraining data that can be generated from simulated environments.
Specifically, the supervised pretraining first constructs a pretraining dataset
@@ 2.2 Supervised pretraining
for $\tau=1,\ldots,t-1$ and $a_{t}^{*}$ is specified by (3). The parameter $\gamma$ and the function $f$ jointly parameterize the distribution $\mathcal{P}_{\gamma,f}$. The parameter $\gamma$ governs the generation of $X_{t}$ and $o_{t}$ just as in (1). The decision function $f$ maps from the history to the action and it governs the generation of $a_{\tau}$’s in the sequence $H_{t}$. In this way, the pretraining data is generated with the following flow:
@@
Based on the dataset $\mathcal{D}_{\text{PT}}$, the supervised pretraining refers to the learning of the decision function $\texttt{TF}_{\theta}$ through minimizing the following empirical loss
@@
where the loss function $l(\cdot,\cdot):\mathcal{A}\times\mathcal{A}\rightarrow\mathbb{R}$ specifies the prediction loss. Thus the key idea of supervised pretraining is to formulate the sequential decision-making task as a sequence modeling task that predicts the optimal action $a_{t}^{(i)*}$ given the history $H_{t}^{(i)}$.
@@
We defer discussions on how various sequential decision-making problems can be formulated under the general setup above and how the pretraining loss is related to the regret to Section B.
3 Pretraining and Generalization
We first point out a subtle point of the supervised pretraining setting presented in the previous section. Unlike supervised learning, the expected loss (8) may not be equal to the expectation of the empirical loss (6). As noted earlier, this is because in the sequence $H_{t}$, the action $a_{t}=\texttt{TF}_{{\hat{\theta}}}(H_{t})$ as in (7) during the test phase as opposed to that $a_{t}=f(H_{t})$ as in (5) for the pertraining data $\mathcal{D}_{\text{PT}}$. Taking expectation with respect to $\gamma$ in (8), we can define the expected loss as
@@ 3 Pretraining and Generalization
Here $f$ is the pre-specified decision function used in generating the pretraining data $\mathcal{D}_{\text{PT}}$. We note that the discrepancy between $L(\texttt{TF}_{\theta})$ and $L_{f}(\texttt{TF}_{\theta})$ is caused by the difference in the generations of the samples $(H_{t},a_{t}^{*})$, namely, $\mathcal{P}_{\gamma,\texttt{TF}_{\theta}}$ versus $\mathcal{P}_{\gamma,f}$; which essentially reduces to how the actions $a_{t}$’s are generated in $H_{t}$. In the training phase, this is generated based on $f$, while measuring the test loss requires that the actions to be generated are based on the transformer $\texttt{TF}_{\theta}.$ This discrepancy is somewhat inevitable because when generating the training data, there is no way we can know the final parameter $\hat{\theta}.$ Consequently, there is no direct guarantee on the loss for $L(\texttt{TF}_{\hat{\theta}};\gamma)$ or $L(\texttt{TF}_{\hat{\theta}})$ with $\hat{\theta}$ being the final learned parameter. This out-of-distribution (OOD) issue motivates the design of our algorithm.
Algorithm 1 Supervised pretraining transformers for sequential decision making
1:Iterations $M$, early training phase $M_{0}\in[1,M]$, number of training sequences per iteration $n$, ratio $\kappa\in[0,1]$, pretraining decision function $f$, prior distribution $\mathcal{P}_{\gamma}$, initial parameter $\theta_{0}$
@@
$L(\texttt{TF}_{\hat{\theta}})$ $\leq\hat{L}(\texttt{TF}_{\hat{\theta}})+\sqrt{\frac{\mathrm{Complex}(\{\texttt{TF}_{\theta}:\theta\in\Theta\})}{nT}}$
$+O\left(\kappa T\cdot\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[W_{1}\left(\mathcal{P}_{\gamma,f},\mathcal{P}_{\gamma,\texttt{TF}_{\hat{\theta}}}\right)\right]+(1-\kappa)T\cdot\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[W_{1}\left(\mathcal{P}_{\gamma,\texttt{TF}_{\tilde{\theta}}},\mathcal{P}_{\gamma,\texttt{TF}_{\hat{\theta}}}\right)\right]+\sqrt{\frac{\log\left(\frac{1}{\delta}\right)}{nT}}\right)$
where $\mathrm{Complex}(\cdot)$ denotes some complexity measure and $W_{1}(\cdot,\cdot)$ is the Wasserstein-1 distance. The notation $O(\cdot)$ omits the boundedness parameters.
@@ 4 Learned Decision Function as an Algorithm
In the previous section, we discuss the training and generalization aspects of the pretrained transformer. Now, we return to the perspective of sequential decision-making and discuss the properties of the decision function when it is used as an algorithm.
@@
The following proposition relates $\texttt{Alg}^{*}(H)$ with the pretraining objective which justifies why it is called the optimal decision function. The loss function $L_{f}(\texttt{Alg})$ is defined as (9); recall that the decision function $f$ is used to generate the actions in $H_{t}$ and the decision function Alg is used to predict $a_{t}^{*}.$
@@
Figure 2: The pretrained transformer $\texttt{TF}_{\hat{\theta}}$ well matches the optimal decision function $\texttt{Alg}^{*}$. For both figures, we plot decision trials for $\texttt{Alg}^{*}$ and $\texttt{TF}_{\hat{\theta}}$. The optimal actions change over time because $X_{t}$’s are different for different time $t$. The experiment setup is deferred to Appendix E.4.
Proposition 4.1 states that the function $\texttt{Alg}^{*}$ is one minimizer of the expected loss under any distribution $\mathcal{P}_{\gamma,f}.$ To see this, $\texttt{Alg}^{*}$ is defined in a pointwise manner for every possible $H.$ The probability $\mathcal{P}_{\gamma,f}$ defines a distribution over the space of $H$. The pointwise optimality of $\texttt{Alg}^{*}$ easily induces the optimality of $\texttt{Alg}^{*}$ for any $\mathcal{P}_{\gamma,f}$. Furthermore, this implies that $\texttt{Alg}^{*}$ is also the optimal solution to the expected test loss and the expected training loss. This means that if the transformer class of $\texttt{TF}_{\theta}$ is rich enough to cover $\texttt{Alg}^{*}$, and with an infinite amount of data, the supervised pretraining will result in $\texttt{Alg}^{*}$ at best. Thus we can interpret the property of the supervised pretrained transformer by $\texttt{Alg}^{*}$ as a proxy. Such a Bayes-optimal function and the Bayesian perspective have been discussed under various settings in the in-context learning literature (Xie et al., 2021; Zhang et al., 2023b; Jeon et al., 2024).
@@
Proposition 4.3 states a negative result that the optimal decision function $\texttt{Alg}^{*}$, albeit being optimal in a prediction sense (for predicting the optimal action), does not serve as a good algorithm for sequential decision-making problems. To see the intuition, the decision function is used recursively in the test phase (see (7)). While it gives the best possible prediction, it does not conduct exploration which is critical for the concentration of the posterior measure. This highlights one key difference between the sequential decision-making problem and the regression problem (Garg et al., 2022) when applying the pretrained transformer. For the regression problem, the in-context samples (in analogy context-action-observation tuples in $H_{t}$) are i.i.d. generated, and this results in a convergence/concentration of the posterior to the true model/environment $\gamma$ (Ghosal et al., 2000). Yet for sequential decision-making problems, the dynamics (7) induces a dependency throughout $H_{t}$, and it may result in a non-concentration behavior for the posterior. And this is the key for constructing the linear-regret example in Proposition 4.3 and the examples are also inspired from and share the same spirit as the discussions in (Harrison et al., 2012). Such lack of exploration does not happen for discrete action space (Lee et al., 2024); this is because for discrete action space, the output $\texttt{TF}_{\hat{\theta}}$ gives the posterior distribution as a distribution over the action space. The distribution output will lead to a randomized action and thus automatically perform exploration.
@@
Consider a finite prior $\mathcal{P}_{\gamma}$ supported on $\Gamma=\{\gamma_{1},...,\gamma_{k}\}$. Suppose (i) The transformer $\texttt{TF}_{\hat{\theta}}$ satisfies $\mathbb{E}[r(X_{t},\texttt{Alg}^{*}(H_{t}))|H_{t}]-\mathbb{E}[r(X_{t},\texttt{TF}_{\hat{\theta}}(H_{t}))|H_{t}]\leq\epsilon$ for every $H_{t}$ where the expectation is taken with respect to the possible randomness in $\texttt{Alg}^{*}$ and $\texttt{TF}_{\hat{\theta}}$. (ii) Denote the conditional distribution of the observation $o_{t}$ given the history $H_{t}$ and the environment $\gamma_{i}$ as $\mathbb{P}_{\gamma_{i}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}}).$ Assume there exists $\Delta\geq 0$ such that the KL divergence
@@
$\text{KL}\left(\mathbb{P}_{\gamma_{i}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})\|\mathbb{P}_{\gamma_{j}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})\right)\geq\Delta$
@@
for any $\gamma_{i},\gamma_{j}\in\Gamma$ and $H_{t}$ and the likelihood ratio between $\mathbb{P}_{\gamma_{i}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})$ and $\mathbb{P}_{\gamma_{j}}(\cdot|H_{t},\texttt{TF}_{\hat{\theta}})$ is sub-Gaussian when $\gamma_{i}$ is the sampled environment. Then we have
@@
$\mathrm{Regret}(\texttt{TF}_{\theta};\gamma)\leq O\left(\epsilon T+\frac{\log k}{\Delta}\right),$
@@
where the notation $O(\cdot)$ omits the boundedness parameters and the sub-Gaussian parameter.

### v2

@@ Abstract
In this paper, we consider the supervised pre-trained transformer for a class of sequential decision-making problems. The class of considered problems is a subset of the general formulation of reinforcement learning in that there is no transition probability matrix; though seemingly restrictive, the subset class of problems covers bandits, dynamic pricing, and newsvendor problems as special cases. Such a structure enables the use of optimal actions/decisions in the pre-training phase, and the usage also provides new insights for the training and generalization of the pre-trained transformer. We first note the training of the transformer model can be viewed as a performative prediction problem, and the existing methods and theories largely ignore or cannot resolve an out-of-distribution issue. We propose a natural solution that includes the transformer-generated action sequences in the training procedure, and it enjoys better properties both numerically and theoretically. The availability of the optimal actions in the considered tasks also allows us to analyze the properties of the pre-trained transformer as an algorithm and explains why it may lack exploration and how this can be automatically resolved. Numerically, we categorize the advantages of pre-trained transformers over the structured algorithms such as UCB and Thompson sampling into three cases: (i) it better utilizes the prior knowledge in the pre-training data; (ii) it can elegantly handle the misspecification issue suffered by the structured algorithms; (iii) for short time horizon such as $T\leq 50$, it behaves more greedy and enjoys much better regret than the structured algorithms designed for asymptotic optimality.
@@ 1 Introduction
In recent years, transformer-based models (Vaswani et al., 2017) have achieved great success in a wide range of tasks such as natural language processing (Radford et al., 2019; Bubeck et al., 2023), computer vision (Dosovitskiy et al., 2020), and also reinforcement learning (RL) (Chen et al., 2021). In particular, an offline RL problem can be framed as a sequence prediction problem (Ding et al., 2019) that predicts the optimal/near-optimal/human action based on the observed history. Chen et al. (2021); Janner et al. (2021) pioneer this approach with the Decision Transformer (DT), which leverages the auto-regressive nature of transformers to maximize the likelihood of trajectories in offline datasets. In this paper, we focus on a subset class of reinforcement learning problems, which we call sequential decision-making problems, that has no transition probability matrix compared to a general RL problem. This subset of problems are general enough to capture a wide range of applications including stochastic and linear bandit problems (Lattimore and Szepesvári, 2020), dynamic pricing (Den Boer, 2015; Cohen et al., 2020), and newsvendor problem (Levi et al., 2015; Besbes and Muharremoglu, 2013). Most importantly, the structure brings two benefits: (i) pre-training – the optimal action is either in closed form or efficiently solvable and thus the optimal action can be used as the prediction target in the pre-training data and (ii) understanding – the setting is more aligned with the existing study (Xie et al., 2021; Garg et al., 2022) on the in-context learning ability of the transformer which gives more insight to the working machinery of transformer on decision-making tasks. Moreover, as a side product, our paper extends pre-trained transformers to the application contexts of business, economics, and operations research which are less studied by the existing works.
Figure 1: Comparison between the pre-trained transformer framework and traditional sequential decision-making methods (such as the structured algorithm of UCB and Thompson sampling). For traditional methods, the decision-making agent (or policy) interacts directly with a single environment sampled from the real environment distribution, focusing on exploration and exploitation within that specific environment. In contrast, the pre-trained transformer is trained across multiple environments sampled from a simulator distribution. During pre-training, the transformer collects trajectories and updates its parameters by interacting with diverse environments. Once pre-trained, the transformer functions works as an algorithm and can be applied effectively in the real environment just as the structured algorithms. But unlike the structured algorithms, the pre-trained transformer leverages the huge amount of pre-training data sampled from the simulation environment.
@@
- The pre-training procedure (Section 3): We mathematically formulate the learning setup of supervised pre-trained transformer for sequential decision-making problems, and propose a standardized procedure to generate the pre-training data. In addition, for the training process, we identify an out-of-distribution (OOD) issue largely ignored or unsolved by the existing literature. Specifically, the pre-training data involves two streams of actions – one is the actions in the history (features), and the other is the action to predict (target). Lin et al. (2023) show that when both streams are generated by one structured algorithm (say UCB or Thompson sampling), the pre-training of the transformer will result in an imitation learning procedure that imitates the structured algorithm at best. If one desires to obtain a better algorithm than the structured algorithm, like using the optimal action (Lee et al., 2024) as the target to predict in our case, this will cause the OOD issue (reflected by an $\Theta(\exp T)$ factor in the regret bound of Lin et al. (2023)). To resolve this matter, we propose an algorithmic solution by injecting the transformer-generated action sequence into the pre-training data, and this also draws a connection with the problem of performative prediction (Perdomo et al., 2020).
- Understanding transformers as decision makers (Section 4): The structure of sequential decision-making problems enables the usage of optimal actions in pre-training data, and as noted earlier and will be exemplified throughout our paper, this brings benefits for both training and theoretical understanding. First, we connect the pre-trained transformer with the existing literature studying in-context learning (Xie et al., 2021; Zhang et al., 2023b; Bai et al., 2024) and we establish the pre-trained transformer as a near Bayes-optimal predictor for the optimal action in a Bayesian sense. Next, we show that (i) why the transformer model exhibits a lack of exploration and (ii) how it achieves a paramount or better performance than the structured algorithms. Other than these, an important aspect is that the problem structure of sequential decision-making renders the setup of our pre-trained transformer different from the existing works that also involve interactions with the underlying environment under the general RL setup. For example, the Online Decision Transformer (ODT) (Zheng et al., 2022) lets the transformers interact with the underlying environment during the training phase, but the interactions are limited to a single environment and are designed to explore better trajectories or policies within that environment than the given offline data. We defer a more detailed comparison between the existing usage of pre-trained transformers on general RL problems and our framework to Appendix A.2.
@@
- Numerically, when and why the pre-trained transformer is better than the structured algorithms (Section 5): While there are many structured algorithms such as UCB and Thompson sampling for each of the decision-making problems that the transformer model is applied to, the question is whether and when the transformer can achieve a better performance than these algorithms. We use numerical experiments to attribute the advantage of the pre-trained transformer into three scenarios: (i) it better utilizes the prior knowledge in the pre-training data; (ii) it can elegantly handle the misspecification issue suffered by the structured algorithms; (iii) for short time horizon such as $T\leq 50$ or $100$, it behaves more greedy and enjoys much better regret than the structured algorithms which are designed for asymptotic optimality.
@@ 2 Problem Setup
In this section, we introduce a general formulation for sequential decision-making and describe the setup of supervised pre-training. An agent (decision maker) makes decisions in a sequential environment. The full procedure is described by the sequence
@@
where $\theta$ encapsulates all the parameters of the function. That is, the agent learns from the past interactions with the underlying environment – tuples of $(X_{t},a_{t},o_{t})$’s to optimize future decisions, and this process that transfers the past knowledge to the future is described by the function $\texttt{TF}_{\theta}$. The function $\texttt{TF}_{\theta}$ can take as inputs a variable-length sequence ($H_{t}$ can be of variable length), which is the case for the generative pre-trained transformers (Radford et al., 2019).
@@ 2.1 Environment and performance metrics
2.2 Supervised pre-training
In the classic study of these sequential decision-making problems, the common paradigm is to assume some structure for the underlying environment and/or the reward function, and to design algorithms accordingly. These algorithms are often known as online or reinforcement learning algorithms and are usually hard to combine with prior knowledge such as pre-training data. Comparatively, the supervised pre-training formulates the decision-making problem as a supervised learning prediction task that predicts the optimal or a near-optimal action, and thus it can utilize the vast availability of pre-training data that can be generated from simulated environments.
Specifically, the supervised pre-training first constructs a pre-training dataset
@@ 2.2 Supervised pre-training
for $\tau=1,\ldots,t-1$ and $a_{t}^{*}$ is specified by (3). The parameter $\gamma$ and the function $f$ jointly parameterize the distribution $\mathcal{P}_{\gamma,f}$. The parameter $\gamma$ governs the generation of $X_{t}$ and $o_{t}$ just as in (1). The decision function $f$ maps from the history to the action and it governs the generation of $a_{\tau}$’s in the sequence $H_{t}$. In this way, the pre-training data is generated with the following flow:
@@
Based on the dataset $\mathcal{D}_{\text{PT}}$, the supervised pre-training refers to the learning of the decision function $\texttt{TF}_{\theta}$ through minimizing the following empirical loss
@@
where the loss function $l(\cdot,\cdot):\mathcal{A}\times\mathcal{A}\rightarrow\mathbb{R}$ specifies the prediction loss. Thus the key idea of supervised pre-training is to formulate the sequential decision-making task as a sequence modeling task that predicts the optimal action $a_{t}^{(i)*}$ given the history $H_{t}^{(i)}$.
@@
We defer discussions on how various sequential decision-making problems can be formulated under the general setup above and how the pre-training loss is related to the regret to Section B.
3 Pre-training and Generalization
We first point out a subtle point of the supervised pre-training setting presented in the previous section. Unlike supervised learning, the expected loss (8) may not be equal to the expectation of the empirical loss (6). As noted earlier, this is because in the sequence $H_{t}$, the action $a_{t}=\texttt{TF}_{{\hat{\theta}}}(H_{t})$ as in (7) during the test phase as opposed to that $a_{t}=f(H_{t})$ as in (5) for the pertraining data $\mathcal{D}_{\text{PT}}$. Taking expectation with respect to $\gamma$ in (8), we can define the expected loss as
@@ 3 Pre-training and Generalization
Here $f$ is the pre-specified decision function used in generating the pre-training data $\mathcal{D}_{\text{PT}}$. We note that the discrepancy between $L(\texttt{TF}_{\theta})$ and $L_{f}(\texttt{TF}_{\theta})$ is caused by the difference in the generations of the samples $(H_{t},a_{t}^{*})$, namely, $\mathcal{P}_{\gamma,\texttt{TF}_{\theta}}$ versus $\mathcal{P}_{\gamma,f}$; which essentially reduces to how the actions $a_{t}$’s are generated in $H_{t}$. In the training phase, this is generated based on $f$, while measuring the test loss requires that the actions to be generated are based on the transformer $\texttt{TF}_{\theta}.$ This discrepancy is somewhat inevitable because when generating the training data, there is no way we can know the final parameter $\hat{\theta}.$ Consequently, there is no direct guarantee on the loss for $L(\texttt{TF}_{\hat{\theta}};\gamma)$ or $L(\texttt{TF}_{\hat{\theta}})$ with $\hat{\theta}$ being the final learned parameter. This out-of-distribution (OOD) issue motivates the design of our algorithm.
Algorithm 1 Supervised pre-training transformers for sequential decision making
1:Iterations $M$, early training phase $M_{0}\in[1,M]$, number of training sequences per iteration $n$, ratio $\kappa\in[0,1]$, pre-training decision function $f$, prior distribution $\mathcal{P}_{\gamma}$, initial parameter $\theta_{0}$
@@
$L(\texttt{TF}_{\hat{\theta}})$ $\leq\hat{L}(\texttt{TF}_{\hat{\theta}})+\sqrt{\frac{\mathrm{Comp}(\{\texttt{TF}_{\theta}:\theta\in\Theta\})}{nT}}$
$+O\left(\kappa T\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[W_{1}\left(\mathcal{P}_{\gamma,f},\mathcal{P}_{\gamma,\texttt{TF}_{\hat{\theta}}}\right)\right]+(1-\kappa)T\mathbb{E}_{\gamma\sim\mathcal{P}_{\gamma}}\left[W_{1}\left(\mathcal{P}_{\gamma,\texttt{TF}_{\tilde{\theta}}},\mathcal{P}_{\gamma,\texttt{TF}_{\hat{\theta}}}\right)\right]+\sqrt{\frac{\log\left(\frac{1}{\delta}\right)}{nT}}\right)$
where $\mathrm{Comp}(\cdot)$ denotes some complexity measure and $W_{1}(\cdot,\cdot)$ is the Wasserstein-1 distance. The notation $O(\cdot)$ omits the boundedness parameters.
@@ 4 Learned Decision Function as an Algorithm
In the previous section, we discuss the training and generalization aspects of the pre-trained transformer. Now, we return to the perspective of sequential decision-making and discuss the properties of the decision function when it is used as an algorithm.
@@
The following proposition relates $\texttt{Alg}^{*}(H)$ with the pre-training objective which justifies why it is called the optimal decision function. The loss function $L_{f}(\texttt{Alg})$ is defined as (9); recall that the decision function $f$ is used to generate the actions in $H_{t}$ and the decision function Alg is used to predict $a_{t}^{*}.$
@@
Figure 3: The pre-trained transformer $\texttt{TF}_{\hat{\theta}}$ well matches the optimal decision function $\texttt{Alg}^{*}$. For both figures, we plot decision trials for $\texttt{Alg}^{*}$ and $\texttt{TF}_{\hat{\theta}}$. The optimal actions change over time because $X_{t}$’s are different for different time $t$. The experiment setup is deferred to Appendix E.4.
Proposition 4.1 states that the function $\texttt{Alg}^{*}$ is one minimizer of the expected loss under any distribution $\mathcal{P}_{\gamma,f}.$ To see this, $\texttt{Alg}^{*}$ is defined in a pointwise manner for every possible $H.$ The probability $\mathcal{P}_{\gamma,f}$ defines a distribution over the space of $H$. The pointwise optimality of $\texttt{Alg}^{*}$ easily induces the optimality of $\texttt{Alg}^{*}$ for any $\mathcal{P}_{\gamma,f}$. Furthermore, this implies that $\texttt{Alg}^{*}$ is also the optimal solution to the expected test loss and the expected training loss. This means that if the transformer class of $\texttt{TF}_{\theta}$ is rich enough to cover $\texttt{Alg}^{*}$, and with an infinite amount of data, the supervised pre-training will result in $\texttt{Alg}^{*}$ at best. Thus we can interpret the property of the supervised pre-trained transformer by $\texttt{Alg}^{*}$ as a proxy. Such a Bayes-optimal function and the Bayesian perspective have been discussed under various settings in the in-context learning literature (Xie et al., 2021; Zhang et al., 2023b; Jeon et al., 2024).
@@
Proposition 4.3 states a negative result that the optimal decision function $\texttt{Alg}^{*}$, albeit being optimal in a prediction sense (for predicting the optimal action), does not serve as a good algorithm for sequential decision-making problems. To see the intuition, the decision function is used recursively in the test phase (see (7)). While it gives the best possible prediction, it does not conduct exploration which is critical for the concentration of the posterior measure. This highlights one key difference between the sequential decision-making problem and the regression problem (Garg et al., 2022) when applying the pre-trained transformer. For the regression problem, the in-context samples (in analogy context-action-observation tuples in $H_{t}$) are i.i.d. generated, and this results in a convergence/concentration of the posterior to the true model/environment $\gamma$ (Ghosal et al., 2000). Yet for sequential decision-making problems, the dynamics (7) induces a dependency throughout $H_{t}$, and it may result in a non-concentration behavior for the posterior. And this is the key for constructing the linear-regret example in Proposition 4.3 and the examples are also inspired from and share the same spirit as the discussions in (Harrison et al., 2012). Such lack of exploration does not happen for discrete action space (Lee et al., 2024); this is because for discrete action space, the output $\texttt{TF}_{\hat{\theta}}$ gives the posterior distribution as a distribution over the action space. The distribution output will lead to a randomized action and thus automatically perform exploration.
@@
Consider a finite prior $\mathcal{P}_{\gamma}$ supported on $\Gamma=\{\gamma_{1},...,\gamma_{k}\}$. Suppose there exists constants $\Delta_{\text{Exploit}}$ and $\Delta_{\text{Explore}}$ such that for any $\gamma\in\Gamma$ and $H_{t}$, $\texttt{TF}_{\hat{\theta}}$ and $\texttt{Alg}^{*}(H_{t})$ satisfies
  • $\mathbb{E}[r(X_{t},\texttt{Alg}^{*}(H_{t}))-r(X_{t},\texttt{TF}_{\hat{\theta}}(H_{t}))|H_{t}]\leq\Delta_{\text{Exploit}}$, where the expectation is taken with respect to the possible randomness in $\texttt{Alg}^{*}$ and $\texttt{TF}_{\hat{\theta}}$.
  • The KL divergence
@@
    $\text{KL}\left(\mathbb{P}_{\gamma}(\cdot|X_{t},\texttt{TF}_{\hat{\theta}}(H_{t}))\|\mathbb{P}_{\gamma^{\prime}}(\cdot|X_{t},\texttt{TF}_{\hat{\theta}}(H_{t}))\right)\geq\Delta_{\text{Explore}}$
@@
    for any $\gamma^{\prime}\in\Gamma$, and the log-likelihood ratio between $\mathbb{P}_{\gamma}(\cdot|X_{t},\texttt{TF}_{\hat{\theta}}(H_{t}))$ and $\mathbb{P}_{\gamma^{\prime}}(\cdot|X_{t},\texttt{TF}_{\hat{\theta}}(H_{t}))$ is $C_{\sigma^{2}}\Delta_{\text{Explore}}$-sub-Gaussian for any $\gamma^{\prime}\in\Gamma$.
Then under conditions on the boundedness of the reward and environment parameters, we have
@@
$\mathrm{Regret}(\texttt{TF}_{\hat{\theta}};\gamma)\leq O\left(\Delta_{\text{Exploit}}T+\frac{\log k}{\Delta_{\text{Explore}}}\right)$
@@
for any $\gamma\in\Gamma$, where the notation $O(\cdot)$ omits logarithm terms in $T$ and the parameters in the boundedness assumptions.